# Stable Voting and the Splitting of Cycles

### v1

@@ 2 Preliminaries
An ordinal margin matrix is a pair $\mathcal{M}={(X,\succ)}$ of a nonempty finite set $X$ (of alternatives) and a strict weak order (i.e., asymmetric and negatively transitive relation) $\succ$ on $X^{2}$ satisfying ordinal skew-symmetry: for all $a,b,c,d\in X$,
@@ 3 Proof of Conjecture for $\leq 5$ Alternatives
and hence $a_{n}$ has a positive margin over $a_{n+1}$. If the margin of $a_{n}$ over $a_{n+1}$ were smaller than that of $b$ over $a$, we claim $a$ would not be the SSV winner, a contradiction. For by ordinal skew-symmetry, the margin of $a_{n+1}$ over $a_{n}$ would be larger than that of $a$ over $b$, in which case we would have considered the match $a_{n+1}$ vs. $a_{n}$ before the match $a$ vs. $b$ in step 2 of Definition 6. Because $a_{n+1}$ is the SSV winner in $\mathcal{M}_{-a_{n}}$, the match $a$ vs. $b$ would not have been the first match in the list of matches by descending margin such that $a$ wins in $\mathcal{M}_{-b}$.
@@
Our proof does not generalize. It finds cycles $C_{1},...,C_{n}$ with increasingly strong edges, ending with an impossible $C_{n}$. $C_{1}$ always exists, but it is one of $O(2^{n}/n)$ cycles, and some $C_{i}$ may block further induction by including none of the $O(n)$ alternatives who wins when either of two others is removed.
@@ 4 SAT encoding
Given any set $\{x_{1},\dots,x_{k}\}$ of propositional variables, the set $\mathcal{L}(x_{1},\dots,x_{k})$ is the smallest set of strings over $\{x_{1},\dots,x_{k},(,),\neg,\wedge,\vee,\to\}$ such that:
@@
    if $\varphi$ and $\psi$ are in $\mathcal{L}(x_{1},\dots,x_{k})$, so are $\neg\varphi$, $(\varphi\wedge\psi)$, $(\varphi\vee\psi)$, and $(\varphi\to\psi)$.
We call elements of $\mathcal{L}(x_{1},\dots,x_{k})$ the Boolean formulas that can be constructed from $x_{1},\dots,x_{k}$. Examples include $x_{1}\to\neg x_{2}$}, $x_{3}\wedge(x_{2}\vee x_{4})$, etc. (omitting outermost parentheses since no ambiguity will arise).
@@
A truth-value assignment $v$ for $\{x_{1},\dots,x_{k}\}$ is a function mapping each variable $x_{i}$ to either True or False. Such an assignment extends recursively to an arbitrary formula $\varphi$ in $\mathcal{L}(x_{1},\dots,x_{k})$ as follows:
@@ 4.1 Linear Ordinal Margin Matrices
We begin by encoding the property of being a linear ordinal margin matrix over $n$ alternatives. Without loss of generality we assume the alternative set $X$ is $[n]$, the set of numbers from $1$ to $n$. We begin by introducing Boolean variables $s_{(a,b),(c,d)}$ for all $a,b,c,d\in[n]$ with $a\neq b$, $c\neq d$, and $(a,b)\neq(c,d)$, representing $(a,b)\succ(c,d)$ in the ordinal margin matrix. For all such variables, we add the following constraints to our formula:
@@
Given an assignment $v$, define $\mathcal{M}_{v}=([n],\succ)$ by
$(a,b)\succ(c,d)\mbox{ iff }v(s_{(a,b),(c,d)})=\textsc{True}.$
If $v$ satisfies the constraints above, then $\mathcal{M}_{v}$ is an ordinal margin matrix.
@@ 4.3 Split Cycle
We prove by induction on $k\in\mathbb{N}$ that if $d$ is reachable from $b$ using a path of length $k$ whose edges are stronger than $(a,b)$ in $\mathcal{M}_{v}$, then $v(r_{(a,b),d})=\textsc{True}$. For $k=0$, one of our constraints is that $v(r_{(a,b),b})=\textsc{True}$. Now assuming the claim holds for paths of length $k$, suppose that $d$ is reachable from $b$ using a path of length $k+1$ whose edges are stronger than $(a,b)$ in $\mathcal{M}_{v}$. Call the last edge in the path $(c,d)$. Then $c$ is reachable from $b$ using a path of length $k$ whose edges are stronger than $(a,b)$, so by the inductive hypothesis, $v(r_{(a,b),c})=\textsc{True}$, and by definition of $\mathcal{M}_{v}$, $(c,d)\succ(d,c)$ and $(c,d)\succ(a,b)$ together imply $v(s_{(c,d),(d,c)}\wedge s_{(c,d),(a,b)})=\textsc{True}$. Then since $v$ satisfies (3), $v(r_{(a,b),d})=\textsc{True}$, which completes the proof.∎
@@
Given an ordinal margin matrix $\mathcal{M}=(X,\succ)$, let $v_{\mathcal{M}}$ be an assignment such that:
@@
There exists a linear ordinal margin matrix in which the SSV winner is not an SC winner iff, where $\varphi$ is the conjunction of all constraints introduced above, $SSV_{[n],1}\wedge\neg SC_{1}\wedge\varphi$} is satisfiable.
@@ 4.5 Results
Using the reasoning from the proof of Theorem 9, if 1 is SC-defeated, it must be by $n$. Otherwise, there exists a counterexample to the conjecture of smaller size. To search for a minimal counterexample, we add the constraint $\lnot SC_{1}\land SC_{n}\land s_{(n,1),(1,n)}\wedge\lnot r_{(n,1),n}$, which asserts that 1 is not an SC winner, that $n$ is, and that $n$ SC-defeats 1.
@@
Although our SAT encoding can be formally verified using Lean (Holliday et al. 2024), formal verification does not seem a pressing issue in this case, since we have a human readable proof up to $n=5$, a verified DRAT proof for $n=6$, and satisfiability results for $n>6$, which we verified using the pref_voting package (Holliday and Pacuit 2025, https://pref-voting.readthedocs.io/en/latest/).

### v2

@@ 2 Preliminaries
An ordinal margin matrix is a pair $\mathcal{M}={(X,\succ)}$ of a nonempty finite set $X$ (of alternatives) and a strict weak order (i.e., an asymmetric and negatively transitive relation) $\succ$ on $X^{2}$ satisfying ordinal skew-symmetry: for all $a,b,c,d\in X$,
@@ 3 Proof of Conjecture for $\leq 5$ Alternatives
and hence $a_{n}$ has a positive margin over $a_{n+1}$. Suppose for contradiction that the margin of $a_{n}$ over $a_{n+1}$ is smaller than that of $b$ over $a$. Then by ordinal skew-symmetry, the margin of $a_{n+1}$ over $a_{n}$ is larger than that of $a$ over $b$, in which case we consider the match $a_{n+1}$ vs. $a_{n}$ before the match $a$ vs. $b$ in step 2 of Definition 6. Then because $a_{n+1}$ is the SSV winner in $\mathcal{M}_{-a_{n}}$, the match $a$ vs. $b$ is not the first match in the list of matches by descending margin such that the first candidate wins after removing the second, contradicting our assumption at the beginning of the proof.
@@
Our proof does not generalize. It finds cycles $C_{1},...,C_{n}$ with increasingly strong edges, ending with an impossible $C_{n}$. $C_{1}$ always exists, but $\mathcal{M}$ may have exponentially many cycles, and some $C_{i}$ may block induction by including none of the alternatives that win when either of two others is removed.
@@ 4 SAT Encoding
Given any set $\{x_{1},\dots,x_{k}\}$ of propositional variables, the set $L(x_{1},\dots,x_{k})$ is the smallest set of strings over $\{x_{1},\dots,x_{k},(,),\neg,\wedge,\vee,\to\}$ such that:
@@
    if $\varphi$ and $\psi$ are in $L(x_{1},\dots,x_{k})$, so are $\neg\varphi$, $(\varphi\wedge\psi)$, $(\varphi\vee\psi)$, and $(\varphi\to\psi)$.
We call elements of $L(x_{1},\dots,x_{k})$ the Boolean formulas that can be constructed from $x_{1},\dots,x_{k}$. Examples include $x_{1}\to\neg x_{2}$}, $x_{3}\wedge(x_{2}\vee x_{4})$, etc. (omitting outermost parentheses since no ambiguity will arise).
@@
A truth-value assignment $v$ for $\{x_{1},\dots,x_{k}\}$ is a function mapping each variable $x_{i}$ to either True or False. Such an assignment extends recursively to an arbitrary formula $\varphi$ in $L(x_{1},\dots,x_{k})$ as follows:
@@ 4.1 Linear Ordinal Margin Matrices
We begin by encoding the property of being a linear ordinal margin matrix over $n$ alternatives. Without loss of generality we assume the alternative set $X$ is $[n]$, the set of numbers from $1$ to $n$. We begin by introducing variables $s_{(a,b),(c,d)}$ for all $a,b,c,d\in[n]$ with $a\neq b$, $c\neq d$, and $(a,b)\neq(c,d)$, representing $(a,b)\succ(c,d)$ in the ordinal margin matrix. For all such variables, we add the following constraints to our formula (where $e\neq f$, $(e,f)\neq(a,b)$, and $(e,f)\neq(c,d)$):
@@
Given an assignment $v$, define $\mathcal{M}_{v}=([n],\succ)$ by $(a,b)\succ(c,d)$ iff (i) $s_{(a,b),(c,d)}$ exists and $v(s_{(a,b),(c,d)})=\textsc{True}$ or (ii) $a=b$, $c\neq d$, and $v(s_{(d,c),(c,d)})=\textsc{True}$ or (iii) $a\neq b$, $c=d$, and $v(s_{(a,b),(b,a)})=\textsc{True}$. If $v$ satisfies the constraints above, then $\mathcal{M}_{v}$ is an ordinal margin matrix.
@@ 4.3 Split Cycle
We prove by induction on $k\in\mathbb{N}$ that if $d$ is reachable from $b$ using a path of length $k$ whose edges are stronger than $(a,b)$ in $\mathcal{M}_{v}$, then $v(r_{(a,b),d})=\textsc{True}$. For $k=0$, one of our constraints is that $v(r_{(a,b),b})=\textsc{True}$. Now assuming the claim holds for paths of length $k$, suppose that $d$ is reachable from $b$ using a path of length $k+1$ whose edges are stronger than $(a,b)$ in $\mathcal{M}_{v}$. Call the last edge in the path $(c,d)$. Then $c$ is reachable from $b$ using a path of length $k$ whose edges are stronger than $(a,b)$, so by the inductive hypothesis, $v(r_{(a,b),c})=\textsc{True}$, and by definition of $\mathcal{M}_{v}$, $(c,d)\succ(d,c)$ and $(c,d)\succ(a,b)$ together imply $\hat{v}(s_{(c,d),(d,c)}\wedge s_{(c,d),(a,b)})=\textsc{True}$. Then since $v$ satisfies (3), $v(r_{(a,b),d})=\textsc{True}$, which completes the proof.∎
@@
Given an ordinal margin matrix $\mathcal{M}=(X,\succ)$, let $v_{\mathcal{M}}$ be an assignment such that for all $s_{(a,b),(c,d)}$:
@@
Where $\varphi$ is the conjunction of all constraints introduced above, there exists a linear ordinal margin matrix in which the SSV winner is not an SC winner iff $SSV_{[n],1}\wedge\neg SC_{1}\wedge\varphi$} is satisfiable.
@@ 4.5 Results
Using the reasoning from the proof of Theorem 9, if 1 is SC-defeated, it must be by $n$. Otherwise, there exists a counterexample to the conjecture of smaller size. To search for a minimal counterexample, we add the constraint $SC_{n}\land s_{(n,1),(1,n)}\wedge\lnot r_{(n,1),n}$ to the formula in Proposition 16.
@@
Although our SAT encoding can be formally verified using Lean (Holliday et al. 2024), formal verification does not seem a pressing issue in this case, since we have a human-readable proof up to $n=5$, a verified DRAT proof for $n=6$, and satisfiability results for $n>6$, which we verified using the pref_voting package (Holliday and Pacuit 2025, https://pref-voting.readthedocs.io/en/latest/).